Determine the annual net cost savings if the robot is purchased. (Do not include the $412,000 inventory reduction or the salvage value in this computation.Saxon Products, Inc., is investigating the purchase of a robot for use on the companys assembly line. Selected data relatingEXHIBIT 13 B-1 Present Value of $1:17 Periods 4% 5% 6% 7% 8% 9% 10% 11% 12% 13% 14% 15% 16% 17% 18% 19% 20% 21% 22% 23% 24% 2EXHIBIT 13B-2 Present Value of an Annuity of $1 in Periods 4% 5% 6% 7% 8% 9% 10% 11% 12% 13% 14% 15% 16% 17% 18% 19% 20% 21%
Answer #1

1.Annual net cost savings:

Savings in inventory carrying cost = $222,000

Savings in labor cost = 23,000*18 = $414,000

Less: Increase in power and maintenance costs = $42,000

Annual net cost savings = $594,000

2.NPV = Present value of cash inflows – present value of cash outflows

= - 1,550,000-405,000 + 594,000*PVAF(23%, 5 years) + 71,000*PVF(23%, 5 years)+412,000*PVF(23%, 1 year) – 412,000*PVF(23%, 5 years)

= -1,955,000 + 594,000*2.803 +71,000*0.355 +412,000*0.813 – 412,000*0.355

= -$76,117

2-b No, since NPV is negative

3-a NPV = -1,955,000 – 87,000 + (18,500*18+222,000-42,000)*2.803 + 71,000*0.355 + 412,000*0.813 – 412,000*0.355

= -$390,160

3-b NO

4-a Increased productivity

Better results

4-b Annual value required = 390,160/2.803

= $139,193.72